# High-Performance Multi-Qubit System with Double-Transmon Couplers towards Scalable Superconducting Quantum Computers

### v1

@@ Abstract
Tunable couplers in superconducting quantum computers have enabled fast and accurate two-qubit gates, with reported high fidelities over 0.99 in various architectures and gate implementation schemes. However, there are few tunable couplers whose performance in multi-qubit systems is clarified, except for the most widely used one: single-transmon coupler (STC). Achieving similar accuracy to isolated two-qubit systems remains challenging due to various undesirable couplings but is necessary for scalability. In this work, we numerically analyze a system of three fixed-frequency qubits coupled via two double-transmon couplers (DTCs) where nearest-neighbor qubits are highly detuned and also next nearest-neighbor ones are nearly resonant. The DTC is a recently proposed tunable coupler, which consists of two fixed-frequency transmons coupled through a common loop with an additional Josephson junction. We find that the DTC can not only reduce undesired residual couplings sufficiently, as well as in isolated two-qubits systems, but also enables implementations of 30-ns CZ gates and 10-ns $\pi/2$ pulses with fidelities of 0.9999 or higher. For comparison, we also investigate the system where the DTCs are replaced by the STCs. The results show that the DTC outperforms the STC in terms of both residual coupling suppression and gate accuracy in the above systems. From these results, we expect that the DTC architecture is promising for realizing high-performance, scalable superconducting quantum computers.
@@ I Introduction
Figure 1: Layout of a typical multiqubit system with the DTCs. Squares and circles represent qubits and coupler transmons, respectively. The colors of qubits (red or blue) indicate which frequency bands (low or high) they belong to. DTCs are represented by two coupler transmons connected via a thin line. The highlighted part, the circuit diagram of which is shown in Fig. 2(a), corresponds to the three-qubit system that we study in this work.
@@
To overcome this disadvantage of the STC, a new kind of tunable coupler named the double-transmon coupler (DTC) has recently been proposedgoto2022DTC . The DTC consists of two fixed-frequency transmons coupled through a common loop with an adittional Josephson junctiongoto2022DTC ; KG2023 ; Airforce2023 . This coupler can make the residual $ZZ$ coupling strength completely zero or negligibly small, typically less than 5 kHz, even for highly detuned qubitsgoto2022DTC ; KG2023 . Also, numerical simulations using superconducting-circuit models have shown that a fast CZ gate using a tunable longitudinal couplinggoto2022DTC ; KG2023 and a fast $\sqrt{\rm iSWAP}$ gate using a parametric transvese couplingKG2023 with fidelities over 0.9999 can be implemented even for highly detuned fixed-frequency qubits. However, all studies of the DTC reported so far have been done only in isolated two-qubit systemsgoto2022DTC ; KG2023 ; Airforce2023 and therefore it has not been clear whether the above abilities of the DTC can be maintained in multi-qubit systems like Fig. 1. In other words, the scalability of the DTC architecture has not been understood well, unlike the other architectures including the STCspremacy1 ; spremacy2 ; TheoriticalSTCcrosstalk2022 ; YanSTCCZ ; sycamore2 ; chinese2 ; crosstalk1 ; crosstalk2 ; crosstalk3 ; crosstalk4 ; crosstalk5 . Since there are additional error sources in multi-qubit systems, such as unwanted interactions between nonadjacent qubits and between couplers, the detailed study on multi-qubit systems with multiple DTCs is highly desirable.
@@
In this paper, we numerically study the system with three qubits coupled via two DTCs as a minimal model for the above purpose (see the highlighted part in Fig. 1). As a result, we find that $ZZ$ coupling between nearest-neighbor (NN) qubits can be reduced as in isolated two-qubit systems. Furthermore, we show that the $ZZ$ coupling between next-nearest-neighbor (NNN) qubits and $ZZZ$ coupling, which do not exist in isolated two-qubit systems, can also be suppressed down to about 1 kHz. We also numerically demonstrate that 30-ns CZ gates and 10-ns $\pi$/2 pulses can be implemented with fidelities of 0.9999 or higher. These results imply that the DTC works well even in multi-qubit systems. Moreover, as a comparison of the above results, we also evaluate the performance of the system where the two DTCs are replaced by the two STCs. The parameters of STCs have been set based on the experiment reporting a high-performance CZ gateETH . The results show that the DTC outperforms the STC in terms of both residual coupling suppression and gate accuracy in the above systems. We find that the difference is mainly due to the difference in parasitic coupling between the NNN qubits.
@@
This paper is organized as follows. In Sec. II, we introduce a theoretical model of the above three-qubit system in Fig. 1 and show the present parameter setting. In Sec. III.1, we show the numerical results of the $ZZ$ and $ZZZ$ couplings and also two-qubit and single-qubit gates (CZ gates and $\pi$/2 pulses, respectively). For the evaluation of gate performance, we use the average gate fidelities fidelity ; kueng2016comparing for the three-qubit system, which can include the effects of spectator errorsTheoriticalSTCcrosstalk2022 ; spectator1 ; spectator2 ; spectator3 ; spectator4 ; spectator5 ; spectator6 ; spectator7 . In Sec. IV, for comparison, we also analyze the system where the DTCs are replaced by the STCs. In Sec. V, we discuss the difference between the results of the DTCs and the STCs. Finally, we summarize our work in Sec. VI.
@@ II.1 Circuit
We consider the three-qubit system shown by the highlighted part in Fig. 1. A corresponding circuit diagram is shown in Fig. 2(a). We also introduce the isolated two-qubit subsystems L and R shown in Figs. 2(b) and 2(c), respectively, corresponding to the left and right sides of the three-qubit system.
@@
II.2 Hamiltonian
We assume $V_{i}=0$ except for single-qubit gates in Sec. III.2.2. The Hamiltonian of this system is then written as follows (see Appendix A for the derivation):
@@ II.2 Hamiltonian
The Hamiltonian in Eq. (1) is represented by a $(2N+1)^{7}\times(2N+1)^{7}$ matrix, where $N$ is a cutoff for the Cooper-pair number (see Appendix B). In this work, we choose $N=10$ for sufficient convergence of energies. Note that calculations of this size are too heavy to perform in a naive manner. To reduce the computational cost, we use the dimension reduction technique explained in Appendix C. We have confirmed that calculated energies and gate fidelities converge with errors of the orders of sub-kHz and $10^{-5}$, respectively. Therefore, in this work we evaluate residual couplings up to 1-kHz rounding off the sub-kHz fractions and the gate fidelities up to 4-digit precision rounding off the fifth decimal place.
@@ II.3 Parameter setting
Figure 3: Energy levels and $ZZ$ coupling strengths of the systems in Fig. 2. (a) and (b) show eigenfrequencies $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5},C_{6},C_{7}}$ as functions of $\Theta_{\rm ex,L}$ and $\Theta_{\rm ex,R}$, respectively. The qubit frequencies of the three-qubit system, $\omega_{Q_{1},Q_{2},Q_{3}}$ $(Q_{i}\in\{0,1\})$, are highlighted by being colored and bold, and the ones of the two-qubit system, $\omega_{Q_{i},Q_{j}}^{\mu}$ ($Q_{i},Q_{j}\in\{0,1\}$ and $\mu\in\{{\rm L,R}\})$ (see Appendix D), are shown by scatter plots. Eigenfrequencies $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5},C_{6},C_{7}}$ other than the above are represented by thin curves. (c) and (d) represent $ZZ$ and $ZZZ$ couplings as functions of $\Theta_{\rm ex,L}$ and $\Theta_{\rm ex,R}$, respectively. Solid curves represent the ones of the three-qubit system and scatter plot represent $\zeta_{12(23)}$ in the system L(R). The orange (blue) dotted horizontal line is $\zeta_{\rm Id,23(12)}$ in the system R(L). Note that $\Theta_{\rm ex,R}$ in the left column and $\Theta_{\rm ex,L}$ in the right column are fixed to $\Theta_{\rm Id,R}$ and $\Theta_{\rm Id,L}$, respectively.
@@
Here, we explain our parameter setting. (1) We set the detunings of the NN qubits, $(\rm{Q_{1}},\rm{Q_{2}})$ and $(\rm{Q_{2}},\rm{Q_{3}})$, to be larger than the absolute values of the qubit anharmonicities $W_{ii}$ $(i\in\{1,2,3\})$. Such a parameter regime, called the highly-detuned regime or out-of-straddling regimeIBMSTC ; goto2022DTC , is preferable to suppress microwave crosstalk between the NN qubits, compared to the nearly resonant or in-the-straddling regimeTheoriticalSTCcrosstalk2022 . (2) From the perspective of suppressing microwave crosstalk, it is desirable to have a larger detuning between the NNN qubits $(\rm{Q_{1}},\rm{Q_{3}})$ as well. However, alternating red-band (lower-frequency) and blue-band (higher-frequency) qubits to keep the NN qubits highly detuned as shown in Fig. 1, the NNN qubits belong to the same-frequency band and thus must nearly resonate. Thus, in this work, we set the detuning between the NNN qubits to a sufficiently small value (10 MHz) compared to the anharmonicity. We will show that even with such a small detuning, crosstalk between the NNN qubits is negligible in this system, but it is not the case for the STCs. The origin of this peformance difference will be discussed in Sec.  V. (3) In actual circuits, there are parasitic capacitances, e.g. $C_{12}$, $C_{13}$, and $C_{56}$, which are not shown in Fig. 2. These capasitances may degrade the performance of the DTCs. Therefore, they should not be ignored. Thus, we set them to feasible values estimated by electromagnetic simulator ANSYS Q3DAnsys . (4) We set $\omega_{i}$ ($i\in\{4,5,6,7\}$) and $r_{Ji}$ $(i\in\{8,9\})$ such that they lead to small residual $ZZ$ couplings and a high-performance adiabatic CZ gate operation.
@@ III Numerical results
III.1 Idling point, $ZZ$ couplings, and $ZZZ$ coupling
@@
Table 3: Average gate fidelities implemented by optimized pulses in the DTC architecture.
$\bar{F}_{CZ,12}$ $\bar{F}_{CZ,23}$ $\bar{F}_{\frac{\pi}{2},1}$ $\bar{F}_{\frac{\pi}{2},2}$ $\bar{F}_{\frac{\pi}{2},3}$
Three-qubit system 0.9999 0.9999 1.0000 1.0000 1.0000
Two-qubit system 1.0000 1.0000 ——– ——– ——–
Figure 3(a)[3(b)] shows $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5},C_{6},C_{7}}$ as functions of $\Theta_{\rm L(R)}$ when $\Theta_{\rm ex,R(L)}$ is fixed to $\Theta_{\rm Id,R(L)}\equiv 0.65\pi\ (0.65\pi)$. Here, $\Theta_{\rm Id,R(L)}$ is an idling point of the isolated two-qubit subsystem L(R) where the $ZZ$ coupling between $\rm{Q_{1}}$ and $\rm{Q_{2}}$ ($\rm{Q_{2}}$ and $\rm{Q_{3}})$ is the closest to 0 (see Appendix D for the details). At the end of this subsection, we will explain the validity that the idling point of the three-qubit system can be chosen as $(\Theta_{\rm ex,L},\Theta_{\rm ex,R})=(\Theta_{\rm Id,L},\Theta_{\rm Id,R})$. The qubit frequencies of the three-qubit system, $\omega_{Q_{1},Q_{2},Q_{3}}$ $(Q_{i}\in\{0,1\})$, are highlighted by being colored and bold, and the ones of the two-qubit system, $\omega_{Q_{i},Q_{j}}^{\mu}$ ($Q_{i},Q_{j}\in\{0,1\}$ and $\mu\in\{{\rm L,R}\})$ (see Appendix D), are shown by scatter plots. We find that $\omega_{1,0,0}$, $\omega_{0,1,0}$, $\omega_{0,0,1}$, $\omega_{1,1,0}$, and $\omega_{0,1,1}$ are in good agreement with the corresponding ones in the two-qubit systems.
@@
$\zeta_{12}$ $=\omega_{1,1,0}-\left[\omega_{1,0,0}+\omega_{0,1,0}\right],$ (9)
$\zeta_{23}$ $=\omega_{0,1,1}-\left[\omega_{0,1,0}+\omega_{0,0,1}\right].$ (10)
We use the same notation , $\zeta_{12}$ and $\zeta_{23}$, for the corresponding $ZZ$ couplings in systems L and R, respectively (see Appendix D for their definitions). The scatter plot in Fig. 3(c)[3(d)] show $\zeta_{12(23)}$ in system L(R). The above $\Theta_{\rm Id,L(R)}$ has been identified as the point where $|\zeta_{12(23)}|$ takes the smallest value. We find that $\zeta_{12(23)}$ in the three-qubit system as a function of $\Theta_{\rm ex,L(R)}$ [the blue (orange) curve] is in good agreement with the one in system L(R). We also found that $\zeta_{23(12)}$ is almost independent of $\Theta_{\rm ex,L(R)}$ in the three-qubit system and its value is close to the one at the idling point in the isolated two-qubit system R(L) shown by the dotted horizontal line in Fig. 3(c)[3(d)]. This independence is reasonable because $\Theta_{\rm ex,L(R)}$ is the parameter of DTC L(R) only connecting $\rm{Q_{1}}$ and $\rm{Q_{2}}$ ($\rm{Q_{2}}$ and $\rm{Q_{3}}$).
@@
$\zeta_{13}$ $=\omega_{1,0,1}-\left[\omega_{1,0,0}+\omega_{0,0,1}\right],$ (11)
@@
$\ \ \ \ \ \ \ -\left[\zeta_{12}+\zeta_{13}+\zeta_{23}\right].$ (12)
@@
Thus, we conclude that $\zeta_{12}$ and $\zeta_{23}$ are almost unchanged from the ones of the isolated two-qubit systems and also $\zeta_{13}$ and $\zeta_{ZZZ}$ are negligible. Therefore, it is reasonable to take the idling point of the three-qubit system to $(\Theta_{\rm ex,L},\Theta_{\rm ex,R})=(\Theta_{\rm Id,L},\Theta_{\rm Id,R})$, as mentioned above. We represent $\zeta_{12}$, $\zeta_{23}$, $\zeta_{13}$, and $\zeta_{ZZZ}$ at the idling point simply as $\zeta_{{\rm Id},12}$, $\zeta_{{\rm Id},23}$, $\zeta_{{\rm Id,13}}$, and $\zeta_{{\rm Id},ZZZ}$, respectively. Their values are summarized in Table 2.

### v2

# High-performance multiqubit system with double-transmon couplers: Toward scalable superconducting quantum computers

@@ Abstract
Tunable couplers in superconducting quantum computers have enabled fast and accurate two-qubit gates, with reported high fidelities over 99% in various architectures and gate implementation schemes. However, there are few tunable couplers whose performance in multi-qubit systems is clarified, except for the most widely used one: single-transmon coupler (STC). Achieving similar accuracy to isolated two-qubit systems remains challenging due to various undesirable couplings but is necessary for scalability. In this work, we numerically analyze a system of three fixed-frequency qubits coupled via two double-transmon couplers (DTCs) where nearest-neighbor qubits are highly detuned and also next nearest-neighbor ones are nearly resonant. The DTC is a recently proposed tunable coupler, which consists of two fixed-frequency transmons coupled through a common loop with an additional Josephson junction. We find that the DTC can not only reduce undesired residual couplings sufficiently, as well as in isolated two-qubits systems, but also enables implementations of 30-ns CZ gates and individual and simultaneous 10-ns $\pi/2$ pulses with fidelities over 99.99%. For comparison, we also investigate the system where the DTCs are replaced by the STCs. The results show that the DTC outperforms the STC in terms of both residual coupling suppression and gate accuracy in the above systems. From these results, we expect that the DTC architecture is promising for realizing high-performance, scalable superconducting quantum computers.
@@ I Introduction
Figure 1: (a) A typical structure of a multiqubit system with the DTCs. Squares and circles represent qubits and coupler transmons, respectively. The colors of qubits (red or blue) indicate which frequency bands (low or high) they belong to. DTCs are represented by two coupler transmons connected via a thin line. (b) A typical layout of the highlighted part of (a) corresponding to the three-qubit system that we study in this work. Its schematic circuit diagram is shown in Fig. 2(a). Its capacitances estimated by electromagnetic simulator ANSYS Q3DAnsys are summarized in Table 1.
@@
To overcome this disadvantage of the STC, a new kind of tunable coupler named the double-transmon coupler (DTC) has recently been proposedgoto2022DTC , numerically investigatedKG2023 ; Airforce2023 , and experimentally realizedRikenToshiba2024 . The DTC consists of two fixed-frequency transmons coupled through a common loop with an addittional Josephson junctiongoto2022DTC ; KG2023 ; Airforce2023 ; RikenToshiba2024 . It has been numericallygoto2022DTC ; KG2023 and experimentallyRikenToshiba2024 demonstrated that the DTC can make the residual $ZZ$ coupling strength completely zero or negligibly small even for highly detuned qubits, which has not been achieved with STCs. Numerical gate simulations using rigorous superconducting-circuit models without decoherence effect have shown that a fast CZ gate using a tunable longitudinal couplinggoto2022DTC ; KG2023 and a fast $\sqrt{\rm iSWAP}$ gate using a parametric transverse couplingKG2023 with fidelities over 99.99% can be implemented for highly detuned fixed-frequency transmon qubits. Furthermore, in the experimental workRikenToshiba2024 , a CZ-gate fidelity of 99.92$\pm$0.01has been realized stably during 12-hour measurement. This is the highest level of two-qubit gate fidelity among superconducting quantum computers ever reported. It has also been reported in Ref. RikenToshiba2024 that the coherence times of qubits coupled via a DTC are also at the highest level as transmons ($T_{1}=228.6,205.3\ \mu$s and $T_{2}^{\rm E}=358.9,129.8\ \mu$s at the idle point). These results suggest that the degradation of the coherence time due to the noise channels introduced by a DTC may be rather smaller than or at least comparable to that of conventional tunable couplers.
@@
However, all studies of the DTC reported so far have been done only in isolated two-qubit systemsgoto2022DTC ; KG2023 ; Airforce2023 ; RikenToshiba2024 and therefore it has not been clear whether the above abilities of the DTC can be maintained in multi-qubit systems like Fig. 1(a). In other words, the scalability of the DTC architecture has not been understood well, unlike the other architectures including the STCspremacy1 ; spremacy2 ; TheoriticalSTCcrosstalk2022 ; YanSTCCZ ; sycamore2 ; chinese2 ; crosstalk1 ; crosstalk2 ; crosstalk3 ; crosstalk4 ; crosstalk5 . Since there are additional error sources in multi-qubit systems, such as unwanted interactions between nonadjacent qubits, between a qubit and a nonadjacent coupler, and between couplers, the detailed study on multi-qubit systems with multiple DTCs is highly desirable.
@@
In this paper, we numerically study the system with three qubits coupled via two DTCs as a minimal model for the above purpose [see the highlighted part in Fig. 1(a)]. As a result, we find that $ZZ$ coupling between nearest-neighbor (NN) qubits can be reduced as in isolated two-qubit systems. Furthermore, we show that the $ZZ$ coupling between next-nearest-neighbor (NNN) qubits and $ZZZ$ coupling, which do not exist in isolated two-qubit systems, can also be suppressed down to about 1 kHz. We also numerically demonstrate that 30-ns CZ gates and individual and simultaneous 10-ns $\pi$/2 pulses can be implemented with fidelities over 99.99%. These results imply that the DTC works well even in multi-qubit systems. Moreover, as a comparison of the above results, we also evaluate the performance of the system where the two DTCs are replaced by the two STCs. The parameters of STCs have been set based on the experiment reporting a high-performance CZ gateETH . The results show that the DTC outperforms the STC in terms of both residual coupling suppression and gate accuracy in the above systems. Based on our findings, the main reason for the difference lies in the fact that the STC architecture typically exhibits a larger stray coupling between the NNN qubits compared to the DTC architecture (see Sec. V).
@@
This paper is organized as follows. In Sec. II, we introduce a theoretical model of the above three-qubit system in Fig. 1 and show the present parameter setting. In Sec. III.1, we show the numerical results of the $ZZ$ and $ZZZ$ couplings and also two-qubit and single-qubit gates (CZ gates and individual and simultaneous $\pi$/2 pulses, respectively). For the evaluation of gate performance, we use the average gate fidelities fidelity ; kueng2016comparing for the three-qubit system, which can include the effects of spectator errorsTheoriticalSTCcrosstalk2022 ; spectator1 ; spectator2 ; spectator3 ; spectator4 ; spectator5 ; spectator6 ; spectator7 . In Sec. IV, for comparison, we also analyze the system where the DTCs are replaced by the STCs. In Sec. V, we discuss the difference between the results of the DTCs and the STCs. Finally, we summarize our work in Sec. VI.
@@ II.1 Circuit
We consider the three-qubit system shown by the highlighted part in Fig. 1(a). A typical device layout and a schematic circuit diagram are shown in Fig. 1(b) and Fig. 2(a), respectively. We also introduce the isolated two-qubit subsystems L and R shown in Figs. 2(b) and 2(c), respectively, corresponding to the left and right sides of the three-qubit system.
@@
II.2 Hamiltonian of the superconducting circuit model
We assume $V_{i}=0$ except for single-qubit gates in Sec. III.2.2. Then, the Hamiltonian of the rigorous circuit model of this system is written as follows (see Appendix A for the derivation):
@@ II.2 Hamiltonian of the superconducting circuit model
In numerical simulations of superconducting quantum computers, effective models derived by approximating the rigorous circuit ones are widely used yan2018tunable ; PRXSTC ; RiggettiSTC ; STCChinese ; IBMSTC ; YanSTCCZ ; TheoriticalSTCcrosstalk2022 . The reason for this is that these models are intuitive and computationally light. However, these effective models may lead to inaccurate results compared to rigorous ones like the above. Therefore, in this study, we use the above rigorous superconducting circuit model without approximations, focusing on demonstrating the performance in an ideal situation in the absence of decoherence as rigorously as possible. In the case of a two-qubit system, it has already been reported that such simulation results are in excellent agreement with experimental onesRikenToshiba2024 .
The Hamiltonian in Eq. (1) is represented by a $(2N+1)^{7}\times(2N+1)^{7}$ matrix, where $N$ is a cutoff for the Cooper-pair number (see Appendix B). In this work, we choose $N=10$ so that the energies converge sufficiently. The matrix size of the Hamiltonian of our three-qubit system $(1801088541\times 1801088541)$ is 85766121 times larger than the one in the two-qubit systems studed in Refs. goto2022DTC ; KG2023 $(194481\times 194481)$. Note that calculations of this size are too heavy to perform in a naive manner, e.g., by directly using QuTipqutip1 ; qutip2 . Due to these difficulties, there have been no studies that conducted gate simulations for multi-qubit systems coupled via tunable coupler using rigorous circuit models, to the best of our knowledge. We overcome this difficulty introducing the dimension reduction technique in Appendix C. We have confirmed that calculated energies and gate fidelities converge with errors of the orders of sub-kHz and $10^{-5}$, respectively. Therefore, in this work we evaluate residual couplings up to 1-kHz rounding off the sub-kHz fractions and the gate fidelities up to 4-digit precision rounding off the fifth decimal place.
@@ II.3 Parameter setting
Here, we explain our parameter setting. (1) We set the detunings of the NN qubits, $(\rm{Q_{1}},\rm{Q_{2}})$ and $(\rm{Q_{2}},\rm{Q_{3}})$, to be larger than the absolute values of the qubit anharmonicities $W_{ii}$ $(i\in\{1,2,3\})$. Such a parameter regime, called the highly-detuned regime or out-of-straddling regimeIBMSTC ; goto2022DTC , is preferable to suppress microwave crosstalk between the NN qubits, compared to the nearly resonant or in-the-straddling regimeTheoriticalSTCcrosstalk2022 . (2) From the perspective of suppressing microwave crosstalk, it is desirable to have a larger detuning between the NNN qubits $(\rm{Q_{1}},\rm{Q_{3}})$ as well. However, alternating red-band (lower-frequency) and blue-band (higher-frequency) qubits to keep the NN qubits highly detuned as shown in Fig. 1, the NNN qubits belong to the same-frequency band and thus must nearly resonate. Thus, in this work, we set the detuning between the NNN qubits to a sufficiently small value (10 MHz) compared to the anharmonicity. We will show that even with such a small detuning, crosstalk between the NNN qubits is negligible in this system, but it is not the case for the STCs. The origin of this peformance difference will be discussed in Sec.  V. (3) In actual circuits, there are parasitic capacitances, e.g. $C_{12}$, $C_{13}$, and $C_{56}$, which are not shown in Fig. 2. These capasitances may degrade the performance of the DTCs. Therefore, they should not be ignored. Thus, we set all of the capacitances, including the parasitic ones, to correspond to the layout shown in Fig. 1(b). We estimate them by the electromagnetic simulator ANSYS Q3DAnsys . We will also discuss the case where the NNN qubits are not on the horizontal line but on the diagonal line in Appendix I. (4) We set $\omega_{i}$ ($i\in\{4,5,6,7\}$) and $r_{Ji}$ $(i\in\{8,9\})$ such that they lead to small residual $ZZ$ couplings and a high-performance adiabatic CZ gate operation.
@@ III Numerical results
Figure 3: Energy levels and $ZZ$ coupling strengths of the systems in Fig. 2. (a) and (b) show eigenfrequencies $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5},C_{6},C_{7}}$ as functions of $\Theta_{\rm ex,L}$ and $\Theta_{\rm ex,R}$, respectively. The qubit frequencies of the three-qubit system, $\omega_{Q_{1},Q_{2},Q_{3}}$ $(Q_{i}\in\{0,1\})$, are highlighted by being colored and bold, and the ones of the two-qubit system, $\omega_{Q_{i},Q_{j}}^{\mu}$ ($Q_{i},Q_{j}\in\{0,1\}$ and $\mu\in\{{\rm L,R}\})$ (see Appendix D), are shown by colored scatter plots. Eigenfrequencies $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5},C_{6},C_{7}}$ other than the above are represented by black thin curves. (c) and (d) represent $ZZ$ and $ZZZ$ couplings as functions of $\Theta_{\rm ex,L}$ and $\Theta_{\rm ex,R}$, respectively. Solid curves represent the ones of the three-qubit system, and scatter plots represent $\zeta_{12(23)}$ in the system L(R). The orange (blue) dotted horizontal line is $\zeta_{\rm Id,23(12)}$ in the system R(L). Note that $\Theta_{\rm ex,R}$ in the left column and $\Theta_{\rm ex,L}$ in the right column are fixed to $\Theta_{\rm Id,R}$ and $\Theta_{\rm Id,L}$, respectively.
III.1 Idle point, $ZZ$ couplings, and $ZZZ$ coupling
@@
Table 3: Average gate fidelities of 30-ns CZ gates and 10-ns individual and simultaneous $\pi/2$ pulses. They are implemented by optimized pulses in the DTC architecture.
Average gate fidelity (%) $\bar{F}_{CZ,12}$ $\bar{F}_{CZ,23}$ $\bar{F}_{\frac{\pi}{2},\{1\}}$ $\bar{F}_{\frac{\pi}{2},\{2\}}$ $\bar{F}_{\frac{\pi}{2},\{3\}}$ $\bar{F}_{\frac{\pi}{2},\{1,2\}}$ $\bar{F}_{\frac{\pi}{2},\{2,3\}}$ $\bar{F}_{\frac{\pi}{2},\{1,3\}}$ $\bar{F}_{\frac{\pi}{2},\{1,2,3\}}$
Three-qubit system 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00
Two-qubit system 100.00 100.00 ——– ——– ——– ——– ——– ——– ——–
Figure 3(a)[3(b)] shows $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5},C_{6},C_{7}}$ as functions of $\Theta_{\rm L(R)}$ when $\Theta_{\rm ex,R(L)}$ is fixed to $\Theta_{\rm Id,R(L)}\equiv 0.65\pi\ (0.65\pi)$. Here, $\Theta_{\rm Id,R(L)}$ is an idle point of the isolated two-qubit subsystem L(R) where the $ZZ$ coupling between $\rm{Q_{1}}$ and $\rm{Q_{2}}$ ($\rm{Q_{2}}$ and $\rm{Q_{3}})$ is the closest to 0 (see Appendix D for the details). At the end of this subsection, we will explain the validity that the idle point of the three-qubit system can be chosen as $(\Theta_{\rm ex,L},\Theta_{\rm ex,R})=(\Theta_{\rm Id,L},\Theta_{\rm Id,R})$. The qubit frequencies of the three-qubit system, $\omega_{Q_{1},Q_{2},Q_{3}}$ $(Q_{i}\in\{0,1\})$, are highlighted by being colored and bold, and the ones of the two-qubit system, $\omega_{Q_{i},Q_{j}}^{\mu}$ ($Q_{i},Q_{j}\in\{0,1\}$ and $\mu\in\{{\rm L,R}\})$ (see Appendix D), are shown by scatter plots. All of them are needed for calculations of $ZZ$ and $ZZZ$ couplings below. We find that $\omega_{1,0,0}$, $\omega_{0,1,0}$, $\omega_{0,0,1}$, $\omega_{1,1,0}$, and $\omega_{0,1,1}$ are in good agreement with the corresponding ones in the two-qubit systems. The black thin curves are frequencies of the states other than the computational state. They cannot be ignored to account for leakage errors during gate simulations.
@@
$\zeta_{12}$ $=\omega_{1,1,0}-\omega_{1,0,0}-\omega_{0,1,0}+\omega_{0,0,0},$ (7)
$\zeta_{23}$ $=\omega_{0,1,1}-\omega_{0,1,0}-\omega_{0,0,1}+\omega_{0,0,0}.$ (8)
We use the same notation , $\zeta_{12}$ and $\zeta_{23}$, for the corresponding $ZZ$ couplings in systems L and R, respectively (see Appendix D for their definitions). The scatter plot in Fig. 3(c)[3(d)] show $\zeta_{12(23)}$ in system L(R). The above $\Theta_{\rm Id,L(R)}$ has been identified as the point where $|\zeta_{12(23)}|$ takes the smallest value. We find that $\zeta_{12(23)}$ in the three-qubit system as a function of $\Theta_{\rm ex,L(R)}$ [the blue (orange) curve] is in good agreement with the one in system L(R). We also found that $\zeta_{23(12)}$ is almost independent of $\Theta_{\rm ex,L(R)}$ in the three-qubit system and its value is close to the one at the idle point in the isolated two-qubit system R(L) shown by the dotted horizontal line in Fig. 3(c)[3(d)]. This independence is reasonable because $\Theta_{\rm ex,L(R)}$ is the parameter of DTC L(R) only connecting $\rm{Q_{1}}$ and $\rm{Q_{2}}$ ($\rm{Q_{2}}$ and $\rm{Q_{3}}$).
@@
$\zeta_{13}$ $=\omega_{1,0,1}-\omega_{1,0,0}-\omega_{0,0,1}+\omega_{0,0,0},$ (9)
@@
$\ \ \ \ \ \ \ -\left[\zeta_{12}+\zeta_{13}+\zeta_{23}\right]+2\omega_{0,0,0}.$ (10)
@@
Thus, we conclude that $\zeta_{12}$ and $\zeta_{23}$ are almost unchanged from the ones of the isolated two-qubit systems and also $\zeta_{13}$ and $\zeta_{ZZZ}$ are negligible. Therefore, it is reasonable to take the idle point of the three-qubit system to $(\Theta_{\rm ex,L},\Theta_{\rm ex,R})=(\Theta_{\rm Id,L},\Theta_{\rm Id,R})$, as mentioned above. We represent $\zeta_{12}$, $\zeta_{23}$, $\zeta_{13}$, and $\zeta_{ZZZ}$ at the idle point simply as $\zeta_{{\rm Id},12}$, $\zeta_{{\rm Id},23}$, $\zeta_{{\rm Id,13}}$, and $\zeta_{{\rm Id},ZZZ}$, respectively. Their values are summarized in Table 2.